Assume the machine shifts and is filling the vials with a mean amount of 9.96 milligrams and a standard deviation of 0.05 milligram. You select five vials and find the mean amount of compound added.



1
Expert's answer
2021-04-13T09:42:22-0400

"μ = 9.96 \\

σ = 0.05 \\

n = 5"

Assumption: the acceptable range of the sample is within +/-1 of the standard deviation.

"P(μ - σ < X <μ + σ ) = P( \frac{(μ - σ) - μ }{σ/ \sqrt{n}} < \frac{X - μ }{σ/ \sqrt{n}}< \frac{(μ - σ) - μ }{σ/ \sqrt{n}} )"


"P( 9.96 - 0.05 < X < 9.96 + 0.05) = P( \frac{( 9.96 - 0.05 ) - 9.96 }{0.05/ \sqrt{5}} < \frac{X - 9.96 }{0.05/ \sqrt{5}} < \frac {( 9.96 + 0.05 ) - 9.96 }{0.05/ \sqrt{5}}) \\


P( 9.995 < X < 9.965 ) = P( \frac{-0.05}{0.05/ \sqrt{5}} < \frac{X - 9.96 }{0.05/ \sqrt{5}} < \frac{0.05 }{0.05/ \sqrt{5}} ) \\


P( 9.995 < X < 9.965 ) = P( -0.4472 < t < 0.4472 ) \\


df = n-1 = 5-1 = 4 \\


P( 9.995 < X < 9.965 ) = 2 \times P( 0 < t < 0.4472 ) \\


P( 9.995 < X < 9.965 ) = 2 \times 0.1632 \\


P( 9.995 < X < 9.965 ) = 0.3264"